# Asymptotic Properties of a Special Solution to the (3,4) String Equation

### v2

@@ Abstract
We analyze the asymptotic properties a special solution of the $(3,4)$ string equation, which appears in the study of the multicritical quartic 22-matrix model. In particular, we show that in a certain parameter regime, the corresponding $\tau$-function has an asymptotic expansion which is ‘topological’ in nature. Consequently, we show that this solution to the string equation with a specific set of Stokes data exists, at least asymptotically. We also demonstrate that, along specific curves in the parameter space, this $\tau$-function degenerates to the $\tau$-function for a tritronquée solution of Painlevé I (which appears in the critical quartic 11-matrix model), indicating that there is a ‘renormalization group flow’ between these critical points. This confirms a conjecture from [CGM90].
@@
In the above, $U=U(t_{5},t_{2},t_{1})$, $V=V(t_{5},t_{2},t_{1})$ and $^{\prime}=\frac{\partial}{\partial t_{1}}$. The above is an ordinary differential equation depending parametrically on the variables $t_{5},t_{2}$. This equation first appeared in the study of the critical 22-matrix model [CGM90, Dou90]. Essentially, the main message of these works was that the critical partition function for the 22-matrix model under study converges in a multi-scaling limit to a $\tau$-function for (1.1). This $\tau$-function then can be interpreted as a partition function for a theory of topological gravity coupled to the Ising conformal field theory [Kri94, DKK93]. A rigorous proof of this statement was finally made in [DHL25]. However, the properties of this solution (even its existence!) have yet to be studied.
@@
One property of the string equation (1.1) is that it has the Painlevé property: its meromorphic solutions have as possible singularities only movable poles and/or fixed branch points. This property is the defining feature of the original 66 Painlevé equations, and is also shared by solutions to the Painlevé hierarchies. Painlevé I – VI all admit a (nonautonomous) Hamiltonian formulation [Oka80], as well as a representation as the isomonodromic deformations of linear differential equations with rational coefficients [JM81, Fok+06]. The same is true for both the Painlevé I and II hierarchies (cf. [Tak07, MM07, CJM06] and references therein). As a consequence of this fact and the seminal works [JMU81, JM81, JM81a], we can associate to each solution of these equations an isomonodromic $\tau$-function, an object which will play an important role in the present work. Furthermore, Painlevé transcendents appear frequently in the context of random matrix theory. Special solutions to the Painlevé I and II hierarchies appear in the universal expressions for critical eigenvalue correlation kernels in random matrix theory [IB03, CV07a], and the partition functions for random matrix models converge (after appropriate normalization) to $\tau$-functions of Painlevé transcendents (see [BD16], and further conjectures of this claim from physics [Gin+90, FGZ95]). Painlevé equations have also found numerous applications in integrable systems [Dub06, CV07] and combinatorics [BDJ99, BD16], among other areas of mathematics and mathematical physics.
@@
All of the aforementioned examples of Painlevé-type equations are examples of rank-2 isomonodromic systems: they can be realized as isomonodromic deformations of some $2\times 2$ matrix-valued linear differential equation with rational coefficients. Aside from some general theory [JMU81, BHH23], not much is known about higher-rank isomonodromic systems, although historically many such systems are of interest to the random matrix theory community. For example, the so-called $(q,p)$ string equations are conjectured to play a role in the classification of universality classes of critical phenomenona in multi-matrix models [Dou90, GM90, Bré+90]. These equations arise from rank $q$ isomonodromic systems, for arbitrary $q\geq 2$. This can be contrasted to the much thinner spectrum of critical phenomena in the 11-matrix model, which are well understood to be indexed by the Painlevé I and II hierarchies, which correspond to the $(2,2g+1)$ and $(2,2g)$ families of string equations, respectively.
@@
The first nontrivial instance of such a higher rank system (i.e., one relevant to random matrix theory [^1] The connection of this equation to random matrix theory is explained in Appendix B., which is not reducible to a rank 22 isomonodromic system) is the one studied in the present work. In other words, (1.1) can be realized as the equation governing the isomonodromic deformations of a rank 33 system. Part of the subject of [Hay24] was to show that although this system is not of rank 22, essentially all of the properties enjoyed by the Painlevé equations are shared by the string equation (1.1). Let us discuss some of these properties here.
@@
Riemann-Hilbert Problem $1.$
@@
This Stokes data is the same as what appears in the multicritical quartic 22-matrix model [DHL25], and so directly applies to this situation. From here on, we will work exclusively with the Riemann-Hilbert Problem 1.1 with Stokes data given by Equation (1).
Remark $1.$
@@
$1.$ Isomonodromic $\tau$-function.
Our main theorems will be stated in terms of the so-called isomonodromic $\tau$-function of Jimbo, Miwa, and Ueno [JMU81]. This is a holomorphic function which is defined in terms of the local data of the RHP $1.$ Usually, the zero locus of a $\tau$-function is precisely the set on which its corresponding Riemann-Hilbert problem is not solvable [Pal99, Mal83]. Furthermore, solutions to the string equation (1.1) can be written as derivatives of the $\tau$-function, and in this sense the $\tau$-function is a fundamental object. However, as pointed out in [Hay24], the definition used in [JMU81] (which we will refer to as the JMU $\tau$ function, and its differential as the JMU differential) does not directly apply to the present situation, as the linear differential equation with rational coefficients associated to RHP 1.1 does not have a diagonalizable leading term. Construction of $\tau$-functions for systems with nondiagonalizable leading terms resonant singularities is a problem that has been studied before [BM05]. In [Hay24], an alternative (albeit less general) formula for such a $\tau$-function was introduced. The advantage of this formula is that it is amenable to steepest descent analysis, as it involves only data derivable from the Riemann-Hilbert problem.
@@
Our first result involves the large-parameter asymptotics of the Riemann-Hilbert problem for $\Psi$: we can say something about the form of the tau function when $t_{5},t_{2},t_{1}$ tend to infinity at a certain rate. This result may be stated explicitly in terms of a particular solution to the following degree 55 equation:
@@
The interpretation of this expansion as one of topological type comes from the relation of this $\tau$-function to the partition function of the multicritical quartic 22-matrix model; this relation is explained further in Appendix B, see Remark B.1. The second main result of this work regards a critical double-scaling limit, and demonstrates the degeneration of this solution to the $(3,4)$ string equation to the tritronquée solution of Painlevé I that appears in the critical 11-matrix model [FIK91, FIK92, DK06, BD16]. These are special solutions to Painlevé I, first studied by Boutroux [Bou13], and are pole-free outside of a sector of angle opening $\frac{2\pi}{5}$ (see also [JK01] for a more modern treatment). As observed by Crnković, Ginsparg and Moore [CGM90], if we make the formal rescaling of variables and take a limit as $T\to+\infty$,  [^2] We put $t_{2}=0$ so that $V\equiv 0$ here for simplicity. We have no reason to doubt that this conjecture (and our proof of it) extend to the $t_{2}\neq 0$ case, if one is willing to work through the technicalities.
@@
where $\varpi_{0}(\eta,0,\nu)$ is as defined in Theorem $1.$ Let us also introduce the notation
@@
  • If AA is an $n\times n$ matrix, and $B$ is an $m\times m$ matrix, we denote by $A\oplus B$ the $(n+m)\times(n+m)$ matrix
@@
We must construct a function $g(\lambda)$, which is defined on a 33-sheeted Riemann surface, and whose restriction $g_{j}(\lambda)$ to sheet $j$ satisfies
@@
and the $g_{j}(\lambda)$ satisfy certain inequalities on the branch cuts, which we will describe in more detail below. The fact that the Riemann surface is 33-sheeted complicates matters. At the present time, the authors do not know an efficient way to treat this problem in general, and so we settle on searching for solutions to the above problem under the assumption that the Riemann surface of $g(\lambda)$ is of genus 0. In this case, this surface can be resolved in terms of rational functions, and we can solve the above problem concretely, albeit with some restrictions on the values of the parameters $\eta,\mu,\nu$. We have the following proposition.
@@
with sheets 11 and 22 glued along $(-\infty,\beta]$, and sheets 22 and 33 glued along $\alpha,\infty)$. Note that, for any values of $(\eta,\mu,\nu)\in D$, we have the inequalities
@@
Since this cubic has 33 real solutions, and so we may use the classical trigonometric Viète formulae to represent these roots:
@@
and so $\text{Re }(g_{3}-g_{2})(\lambda)=0$ for $z\in(\alpha,\infty)$. By the definition of sheets 22 and 33,
@@
$3.$ Proof of Theorems 1.1 & 1.2.
With the steepest descent analysis for $\bf Z$} completed, we can now proceed to the proof of Theorem $1.$
@@
  • If NN is odd and $\ell\in\operatorname{\mathbb{Z}}_{+}$, then either NN-$\ell$ is odd or $\ell$ is odd,
  • If NN is odd, $\ell,j\in\operatorname{\mathbb{Z}}_{+}$, then one of NN-$\ell$, $\ell-j$, $j$ is odd.
@@
where $c:=\left(5s/6\right)^{1/5}>0$. Clearly, the above has determinant 11, and furthermore is analytic in $\operatorname{\mathbb{C}}\setminus D_{\hbar}$. Finally, we define the modified global parametrix
@@
In summary, we have studied some asymptotic properties of a special solution to the $(3,4)$ string equation (1.1), which appears in the study of the multicritical quartic 22-matrix model. In particular, we were able to show that this solution admits a “topological” expansion, and that the $\tau$-function for this solution has limits to the Painlevé I $\tau$-function, confirming a conjecture from [CGM90].
@@
    The Stokes manifold for the symmetric solutions of the RHP 1.1 (i.e., solutions with corresponding Stokes parameters satisfying $s_{k}=-s_{k+8}$, k=7,,1k=-7,...,-1) contains the following 77 planes, which we parameterize in terms of the remaining 77 Stokes parameters $(s_{1},....,s_{7})$:
@@
    These planes intersect pairwise: $\Pi_{k}\cap\Pi_{k\pm 1}\neq\emptyset$ for $k\in\operatorname{\mathbb{Z}}_{7}$, and otherwise are completely disjoint. As these planes are embedded in $\operatorname{\mathbb{C}}^{7}$, their intersection is a point in each case. Note that the Stokes parameters corresponding to the model studied in the present work (1) lies at the intersection of $\Pi_{0}$ and $\Pi_{1}$. This is analogous to the way the tritronquée solutions to Painlevé I (which lie at the intersection points of pairs of the 55 lines defining the tronquée solutions in the PI Stokes manifold, cf. [KK93]) appear in the 11-matrix model. In particular, the set of solutions to the string equation lying on the plane $\Pi_{1}$ with parameter $y=0$ can be studied using the same techniques in this work, with little modification: one only has to choose the relevant corresponding Stokes data in the local model problems accordingly. The result is that our main theorems hold for each of these solutions, with the word tritronquée solution replaced with tronquée solution of Painlevé I. The role of the Stokes parameter $y$ is unclear. Finding an asymptotic formula for solutions to the string equation with generic Stokes data on the plane $\Pi_{1}$ is still in general an open problem.
@@
Interpretation of the above formal asymptotics. Within the context of the 22-matrix model, the above expansion has the following formal interpretation. Recall that the free energy $\mathcal{F}(\tau,t,H)$ of the quartic 22-matrix model:
@@
The Stokes manifold for this Riemann-Hilbert problem contains 55 lines, which can be found by solving Equation (D.3) under the constraint that one of the parameters $\mathfrak{s}_{k}$ is zero. These lines should correspond to the tronquée solutions of Painlevé I. The Stokes data relevant to us is

### v3

@@ Abstract
We analyze the asymptotic properties a special solution of the $(3,4)$ string equation, which appears in the study of the multicritical quartic $2$-matrix model. In particular, we show that in a certain parameter regime, the corresponding $\tau$-function has an asymptotic expansion which is ‘topological’ in nature. Consequently, we show that this solution to the string equation with a specific set of Stokes data exists, at least asymptotically. We also demonstrate that, along specific curves in the parameter space, this $\tau$-function degenerates to the $\tau$-function for a tritronquée solution of Painlevé I (which appears in the critical quartic $1$-matrix model), indicating that there is a ‘renormalization group flow’ between these critical points. This confirms a conjecture from [CGM90].
@@
In the above, $U=U(t_{5},t_{2},t_{1})$, $V=V(t_{5},t_{2},t_{1})$ and $^{\prime}=\frac{\partial}{\partial t_{1}}$. The above is an ordinary differential equation depending parametrically on the variables $t_{5},t_{2}$. This equation first appeared in the study of the critical $2$-matrix model [CGM90, Dou90]. Essentially, the main message of these works was that the critical partition function for the $2$-matrix model under study converges in a multi-scaling limit to a $\tau$-function for (1.1). This $\tau$-function then can be interpreted as a partition function for a theory of topological gravity coupled to the Ising conformal field theory [Kri94, DKK93]. A rigorous proof of this statement was finally made in [DHL25]. However, the properties of this solution (even its existence!) have yet to be studied.
@@
One property of the string equation (1.1) is that it has the Painlevé property: its meromorphic solutions have as possible singularities only movable poles and/or fixed branch points. This property is the defining feature of the original $6$ Painlevé equations, and is also shared by solutions to the Painlevé hierarchies. Painlevé I – VI all admit a (nonautonomous) Hamiltonian formulation [Oka80], as well as a representation as the isomonodromic deformations of linear differential equations with rational coefficients [JM81, Fok+06]. The same is true for both the Painlevé I and II hierarchies (cf. [Tak07, MM07, CJM06] and references therein). As a consequence of this fact and the seminal works [JMU81, JM81, JM81a], we can associate to each solution of these equations an isomonodromic $\tau$-function, an object which will play an important role in the present work. Furthermore, Painlevé transcendents appear frequently in the context of random matrix theory. Special solutions to the Painlevé I and II hierarchies appear in the universal expressions for critical eigenvalue correlation kernels in random matrix theory [IB03, CV07a], and the partition functions for random matrix models converge (after appropriate normalization) to $\tau$-functions of Painlevé transcendents (see [BD16], and further conjectures of this claim from physics [Gin+90, FGZ95]). Painlevé equations have also found numerous applications in integrable systems [Dub06, CV07] and combinatorics [BDJ99, BD16], among other areas of mathematics and mathematical physics.
@@
All of the aforementioned examples of Painlevé-type equations are examples of rank-2 isomonodromic systems: they can be realized as isomonodromic deformations of some $2\times 2$ matrix-valued linear differential equation with rational coefficients. Aside from some general theory [JMU81, BHH23], not much is known about higher-rank isomonodromic systems, although historically many such systems are of interest to the random matrix theory community. For example, the so-called $(q,p)$ string equations are conjectured to play a role in the classification of universality classes of critical phenomenona in multi-matrix models [Dou90, GM90, Bré+90]. These equations arise from rank $q$ isomonodromic systems, for arbitrary $q\geq 2$. This can be contrasted to the much thinner spectrum of critical phenomena in the $1$-matrix model, which are well understood to be indexed by the Painlevé I and II hierarchies, which correspond to the $(2,2g+1)$ and $(2,2g)$ families of string equations, respectively.
@@
The first nontrivial instance of such a higher rank system (i.e., one relevant to random matrix theory [^1] The connection of this equation to random matrix theory is explained in Appendix B., which is not reducible to a rank $2$ isomonodromic system) is the one studied in the present work. In other words, (1.1) can be realized as the equation governing the isomonodromic deformations of a rank $3$ system. Part of the subject of [Hay24] was to show that although this system is not of rank $2$, essentially all of the properties enjoyed by the Painlevé equations are shared by the string equation (1.1). Let us discuss some of these properties here.
@@
Riemann-Hilbert Problem 1.1.
@@
This Stokes data is the same as what appears in the multicritical quartic $2$-matrix model [DHL25], and so directly applies to this situation. From here on, we will work exclusively with the Riemann-Hilbert Problem 1.1 with Stokes data given by Equation (1).
Remark 1.1.
@@
1.1. Isomonodromic $\tau$-function.
Our main theorems will be stated in terms of the so-called isomonodromic $\tau$-function of Jimbo, Miwa, and Ueno [JMU81]. This is a holomorphic function which is defined in terms of the local data of the RHP 1.1. Usually, the zero locus of a $\tau$-function is precisely the set on which its corresponding Riemann-Hilbert problem is not solvable [Pal99, Mal83]. Furthermore, solutions to the string equation (1.1) can be written as derivatives of the $\tau$-function, and in this sense the $\tau$-function is a fundamental object. However, as pointed out in [Hay24], the definition used in [JMU81] (which we will refer to as the JMU $\tau$ function, and its differential as the JMU differential) does not directly apply to the present situation, as the linear differential equation with rational coefficients associated to RHP 1.1 does not have a diagonalizable leading term. Construction of $\tau$-functions for systems with nondiagonalizable leading terms resonant singularities is a problem that has been studied before [BM05]. In [Hay24], an alternative (albeit less general) formula for such a $\tau$-function was introduced. The advantage of this formula is that it is amenable to steepest descent analysis, as it involves only data derivable from the Riemann-Hilbert problem.
@@
Our first result involves the large-parameter asymptotics of the Riemann-Hilbert problem for $\Psi$: we can say something about the form of the tau function when $t_{5},t_{2},t_{1}$ tend to infinity at a certain rate. This result may be stated explicitly in terms of a particular solution to the following degree $5$ equation:
@@
The interpretation of this expansion as one of topological type comes from the relation of this $\tau$-function to the partition function of the multicritical quartic $2$-matrix model; this relation is explained further in Appendix B, see Remark B.1. The second main result of this work regards a critical double-scaling limit, and demonstrates the degeneration of this solution to the $(3,4)$ string equation to the tritronquée solution of Painlevé I that appears in the critical $1$-matrix model [FIK91, FIK92, DK06, BD16]. These are special solutions to Painlevé I, first studied by Boutroux [Bou13], and are pole-free outside of a sector of angle opening $\frac{2\pi}{5}$ (see also [JK01] for a more modern treatment). As observed by Crnković, Ginsparg and Moore [CGM90], if we make the formal rescaling of variables and take a limit as $T\to+\infty$,  [^2] We put $t_{2}=0$ so that $V\equiv 0$ here for simplicity. We have no reason to doubt that this conjecture (and our proof of it) extend to the $t_{2}\neq 0$ case, if one is willing to work through the technicalities.
@@
where $\varpi_{0}(\eta,0,\nu)$ is as defined in Theorem 1.1. Let us also introduce the notation
@@
  • If $A$ is an $n\times n$ matrix, and $B$ is an $m\times m$ matrix, we denote by $A\oplus B$ the $(n+m)\times(n+m)$ matrix
@@
We must construct a function $g(\lambda)$, which is defined on a $3$-sheeted Riemann surface, and whose restriction $g_{j}(\lambda)$ to sheet $j$ satisfies
@@
and the $g_{j}(\lambda)$ satisfy certain inequalities on the branch cuts, which we will describe in more detail below. The fact that the Riemann surface is $3$-sheeted complicates matters. At the present time, the authors do not know an efficient way to treat this problem in general, and so we settle on searching for solutions to the above problem under the assumption that the Riemann surface of $g(\lambda)$ is of genus 0. In this case, this surface can be resolved in terms of rational functions, and we can solve the above problem concretely, albeit with some restrictions on the values of the parameters $\eta,\mu,\nu$. We have the following proposition.
@@
with sheets $1$ and $2$ glued along $(-\infty,\beta]$, and sheets $2$ and $3$ glued along $\alpha,\infty)$. Note that, for any values of $(\eta,\mu,\nu)\in D$, we have the inequalities
@@
Since this cubic has $3$ real solutions, and so we may use the classical trigonometric Viète formulae to represent these roots:
@@
and so $\text{Re }(g_{3}-g_{2})(\lambda)=0$ for $z\in(\alpha,\infty)$. By the definition of sheets $2$ and $3$,
@@
3.3. Proof of Theorems 1.1 & 1.2.
With the steepest descent analysis for $\bf Z$} completed, we can now proceed to the proof of Theorem 1.1.
@@
  • If $N$ is odd and $\ell\in\operatorname{\mathbb{Z}}_{+}$, then either NN-$\ell$ is odd or $\ell$ is odd,
  • If $N$ is odd, $\ell,j\in\operatorname{\mathbb{Z}}_{+}$, then one of NN-$\ell$, $\ell-j$, $j$ is odd.
@@
where $c:=\left(5s/6\right)^{1/5}>0$. Clearly, the above has determinant $1$, and furthermore is analytic in $\operatorname{\mathbb{C}}\setminus D_{\hbar}$. Finally, we define the modified global parametrix
@@
In summary, we have studied some asymptotic properties of a special solution to the $(3,4)$ string equation (1.1), which appears in the study of the multicritical quartic $2$-matrix model. In particular, we were able to show that this solution admits a “topological” expansion, and that the $\tau$-function for this solution has limits to the Painlevé I $\tau$-function, confirming a conjecture from [CGM90].
@@
    The Stokes manifold for the symmetric solutions of the RHP 1.1 (i.e., solutions with corresponding Stokes parameters satisfying $s_{k}=-s_{k+8}$, k=7,,1k=-7,...,-1) contains the following $7$ planes, which we parameterize in terms of the remaining $7$ Stokes parameters $(s_{1},....,s_{7})$:
@@
    These planes intersect pairwise: $\Pi_{k}\cap\Pi_{k\pm 1}\neq\emptyset$ for $k\in\operatorname{\mathbb{Z}}_{7}$, and otherwise are completely disjoint. As these planes are embedded in $\operatorname{\mathbb{C}}^{7}$, their intersection is a point in each case. Note that the Stokes parameters corresponding to the model studied in the present work (1) lies at the intersection of $\Pi_{0}$ and $\Pi_{1}$. This is analogous to the way the tritronquée solutions to Painlevé I (which lie at the intersection points of pairs of the $5$ lines defining the tronquée solutions in the PI Stokes manifold, cf. [KK93]) appear in the $1$-matrix model. In particular, the set of solutions to the string equation lying on the plane $\Pi_{1}$ with parameter $y=0$ can be studied using the same techniques in this work, with little modification: one only has to choose the relevant corresponding Stokes data in the local model problems accordingly. The result is that our main theorems hold for each of these solutions, with the word tritronquée solution replaced with tronquée solution of Painlevé I. The role of the Stokes parameter $y$ is unclear. Finding an asymptotic formula for solutions to the string equation with generic Stokes data on the plane $\Pi_{1}$ is still in general an open problem.
@@
Interpretation of the above formal asymptotics. Within the context of the $2$-matrix model, the above expansion has the following formal interpretation. Recall that the free energy $\mathcal{F}(\tau,t,H)$ of the quartic $2$-matrix model:
@@
The Stokes manifold for this Riemann-Hilbert problem contains $5$ lines, which can be found by solving Equation (D.3) under the constraint that one of the parameters $\mathfrak{s}_{k}$ is zero. These lines should correspond to the tronquée solutions of Painlevé I. The Stokes data relevant to us is